OpenStudy (anonymous):

What is the quotient: (–3x3 + 16x2 + 13x – 6) ÷ (3x – 1)?

OpenStudy (anonymous):

wait nvm (3x - 1) * (-x^2) => (-3x^3 + x^2) (–3x3 + 16x2 + 13x – 6) - (-3x^3 + x^2) = (15x^2 + 13x - 6) (3x - 1) * (5x) => (15x^2 - 5x) (15x^2 + 13x - 6) - (15x^2 - 5x) = (18x - 6) (3x - 1) * 6 = 18x - 6 (18x - 6) - (18x - 6) = 0 So, the final result will be => -x^2 + 5x + 6

OpenStudy (radar):